OpenStudy (anonymous):

Which of the following expressions is equivalent to a9 - b9?

OpenStudy (solomonzelman):

you mean \(\Large\color{black}{ a^9-b^9}\) ?

OpenStudy (anonymous):

Which of the following expressions?? The following expression here is of Solomon and not yours. :P

OpenStudy (anonymous):

@nick305 are you given with some answer options or choices?

OpenStudy (solomonzelman):

I am NOT going to expand it. But you can certainly put it in wolframalpha.com

OpenStudy (anonymous):

yeah solomon

OpenStudy (anonymous):

yeah let me give you the answer choices @waterineyes

OpenStudy (anonymous):

A. (a3)3 - (b3)3 B. (a6)3 - (b6)3 C. (a3 - b3)2 D. (a3- b3)3

OpenStudy (anonymous):

Rule of indices: \[(a^m)^n = a^{m \times n}\]

OpenStudy (anonymous):

so i can eliminate c and d @waterineyes

OpenStudy (anonymous):

Yes for sure.

OpenStudy (anonymous):

What else you can eliminate?

OpenStudy (anonymous):

Can you write \(9\) as \(3 \times 3\) ??

OpenStudy (anonymous):

b

OpenStudy (anonymous):

Good.. :)

OpenStudy (anonymous):

thank you @waterineyes

OpenStudy (anonymous):

See, \(a^9\) in other form can be written as: \[a^9 = a^{3 \times 3} = (a^3)^3\] Just using that indices rule. :)

OpenStudy (anonymous):

\(m=3\) and \(n = 3\) also there. you are welcome dear. :)

OpenStudy (anonymous):

lol you the best

OpenStudy (anonymous):

"Best", that I can't say, but yes I am lol.. :P

OpenStudy (anonymous):

For your next post, any, do tag me there, one thing is remained, that I can't do here. :P

OpenStudy (anonymous):

ok thanks